i divided a number by 52.if the quotient was 47 and remainder was 47,the number was.
1
Expert's answer
2013-01-09T09:11:31-0500
Let the required number is x. So x=52*q+rwhere q - quotient, r - remainder
Therefore we get x=52*47+47=(52+1)*47=53*47=2491